# The Role of Privacy Guarantees in Voluntary Donation of Private Health Data for Altruistic Goals

### v2

@@ 5. Concluding Discussion
Our participants reported strong pre-existing privacy expectations, even in the absence of explicit privacy statements. We observe that these expectations vary strongly based on the profit model of the entity. For for-profit entities, where baseline privacy expectations are lower, strong privacy statements can effectively elevate these expectations among users. In contrast, the impact on non-profit organizations is minimal because users already hold high privacy expectations for these entities.
@@
The Halo Effect.The Halo EffectThe Halo Effect. Participants qualitative responses suggest that their strongly positive privacy expectations for non-profit entities are rooted in arguments of morality: non-profits are perceived as more ethical and dedicated to serving the public good because they are not pursuing profit. This findings are consistent with a phenomenon termed “the halo effect”, which describes how an overall positive impression of a particular entity (e.g., the good work that a non-profit does) can lead to ungrounded assumptions about how the organization operates (Nisbett and Wilson, 1977; de Bruin Cardoso et al., 2024). While prior work in the business literature focuses on how the halo effect influences expectations about business practices such as donation management, our work illustrates that the halo effect also influences people’s expectations about the way non-profits handle data privacy.
These positive expectations do not always align with reality. Privacy safeguards are not consistently enforced or widely practiced; for example, many health organizations rely on broad consent frameworks rather than adopting stricter purpose-restricted consent measures (Kaye et al., 2015). This gap between expectations and reality can result in privacy expectation violations, a major source of privacy concerns as explained by the theory of contextual integrity (Nissenbaum, 2009). Indeed, a well-known non-profit mental health helpline shared it’s data with a commercial entity, leading to significant public outcry (Levine, 2022). Future work may seek to explore how to design privacy communications that set appropriate expectations in the presence of such effects, as a prior work has sought to do in other domains (e.g., (Burton et al., 2015)).
@@
Privacy Washing.Privacy WashingPrivacy Washing. Conversely, our participants’ had far lower initial expectations of the privacy practices of for-profit entities due to their perceived singular focus on profit (a “horn” effect). As a result, information about data protection guarantees significantly raised their privacy expectations, and in turn, their willingness to donate their data.
This means that for-profit entities have the opportunity to gain “competitive advantage” (Schaub et al., 2015) through influencing and enhancing user privacy expectations via transparently implementing and effectively communicating about PETs. However, this effect also posses a vulnerability to ”privacy washing” (MainWP, 2023; Cirucci, 2024), in which organizations make vague or exaggerated claims about PETs to falsely raise privacy expectations. Such deceptive practices can mislead individuals into sharing their data, erode trust, and discourage future data donations if individuals later feel deceived.
@@
The Role of Auditing.The Role of AuditingThe Role of Auditing. In theory, auditing could offer verification that combats both the halo effect and privacy washing by grounding assumed or stated privacy guarantees in technical reality. Unfortunately, we find that auditing has a limited effect on elevating privacy expectations among participants, except in scenarios involving purpose use restrictions by for-profit entities. Participants understand that while PETs and auditing processes are in place, they are not infallible, and auditing does not substantially mitigate the practical risks associated with data breaches or unauthorized access.
This perception shows a disconnect between the perspectives of experts and participants, as shown in (Dechand et al., 2019; Oates et al., 2018; Wu and Zappala, 2018). Experts often regard auditing as a fundamental assurance mechanism that reinforces trust and compliance. In contrast, participants are more concerned with the overall protection of their data, regardless of specific technical implementations or compliance processes. As one respondent noted, ”They can guarantee privacy all they want; things still get hacked.” True privacy is doubted to exist (Lerner et al., 2022).
@@
Building Trust.Building TrustBuilding Trust. Our respondent’s prior experiences with data leakage often informed this fatalistic perspective on privacy. Technologies like multi-party computation (MPC), although not yet scalable, have the potential to minimize the impact of a breach by specifically distributing sensitive information across different entities to ensure that no single entity can reconstruct the information on its own. However, deploying such systems requires careful engineering to prevent distributed data breaches on all the entities involved. Existing skepticism about data security means that any failure of these systems could severely damage trust, potentially more so than if simpler, well-understood methods were compromised.
@@
Future Work.Future WorkFuture Work. In this work, we explicitly examined the influence of four privacy guarantees and two auditing guarantees. Future research could broaden this scope by incorporating a more diverse and comprehensive range of PETs, with a particular focus on understanding how multiple guarantees interact in real-world scenarios where they often coexist. Additionally, researchers can also conduct targeted studies to understand the perspectives of diverse demographic groups, including individuals with a strong distrust of data-handling practices. Identifying and understanding these individuals is critical in fostering confidence in PETs and practices.
Our study focused exclusively on individuals’ perspectives regarding formal medical data donation scenarios. However, medical data are also frequently collected through smart devices and donated with a single click via in-app prompts. Prior work emphasizes the role of modality and timing in the design space of privacy information (Schaub et al., 2015). Future research could explore how such factors intersect with biases such as the halo effect to influence privacy expectations, trust, and willingness to share sensitive information.
Finally, prior work (Seberger and Patil, 2021) studying COVID-19 contract tracing apps – many created through a partnership between for-profits and non-profits – found that people assumed these apps would “operate according to the economics of surveillance capitalism…[making it] unreasonable to expect them to trust the system that profits from the sale of their data.” Future work may seek to further interrogate how such partnerships, which are common in the medical domain, further complicate our understanding of people’s privacy expectations and their sensitivity to statements of data protection.
Appendix A Survey
In this section, we report the full text of our questionnaire. We first ask for a participant themselve or their close relatives have been diagnosed with a disease. In the rest of the questionnaire, we will choose at random one disease the participant has selected, if any, or one disease the participant has not selected. This is labeled in Italic as disease. The entity, also in Italic, may be either non-profit entity or for-profit entity, chosen at random.
Egocentricity
  1. (1)
    Have you ever been diagnosed with any of these diseases?
    • Cancer (Breast, Colorectal, Lung, and Prostate)
    • Diabetes
    • Heart Failure
    • Hypertension (High blood pressure)
    • Stroke
    • None of the above
  2. (2)
    Have your partner and/or close relatives (parents, children, siblings) ever been diagnosed with any of these diseases?
    • Cancer (Breast, Colorectal, Lung, and Prostate)
    • Diabetes
    • Heart Failure
    • Hypertension (High blood pressure)
    • Stroke
    • None of the above
Survey Introduction
Next, we present you with a scenario and ask you how you would respond. Please do your best to imagine yourself in this scenario. There is no right or wrong answer to these questions: we are interested to hear how you would feel about being in this scenario and what you would do.
Imagine that an entity entity wants to develop a new treatment for disease. They need medical data from people with and without disease to develop the treatment. They ask you to donate your medical record to help develop the treatment.
Your medical record contains your: (i) personal information, which may include information about your age, weight, gender, race; (ii) medical history, which may include information about allergies, illnesses, surgeries, immunizations, and results of physical exams and tests; and (iii) medical behavior, which may include information about medicines taken and health habits, such as smoking habits, diet and exercise.
Questions
We will now ask you a series of questions about this scenario. For your reference, the scenario text will appear at the top of each page.
  1. (1)
    How would you rate your understanding of the above scenario?
    • Fully understand
    • Mostly understand
    • Partially understand
    • Not understand
  2. (2)
    In this scenario how likely would you be to donate your medical record?
    • Very likely
    • Likely
    • Unlikely
    • Very unlikely
  3. (3)
    Why your are willing (or unwilling) to share your medical record with this entity?
  4. (4)
    Please rate your agreement with the following statement:
    I trust the entity will handle my data as described.
    • Strongly agree
    • Somewhat agree
    • Somewhat disagree
    • Strongly disagree
  5. (5)
    Please explain why you do (or do not) trust that the entity will handle your data as described.
  6. (6)
    In this scenario, how likely do you think it is that each of the following will occur?
      • Very likely
      • Likely
      • Unlikely
      • Very unlikely
    • The donated medical record will be deleted at a set point in time.
      • Very likely
      • Likely
      • Unlikely
      • Very unlikely
      • Very likely
      • Likely
      • Unlikely
      • Very unlikely
      • Very likely
      • Likely
      • Unlikely
      • Very unlikely
      • Very likely
      • Likely
      • Unlikely
      • Very unlikely
      • Very likely
      • Likely
      • Unlikely
      • Very unlikely
    • If I donate my data, I will meet Albert Einstein.
      • Very likely
      • Likely
      • Unlikely
      • Very unlikely
  7. (7)
    History In real life, have you ever donated your medical record?
    • Yes, I have donated my medical record for disease
    • Yes, I have donated my medical record for research of other illnesses.
    • Yes, I have donated my medical record for non-medical purposes.
    • No, I was asked to donate my medical record but I declined to participate.
    • No, I have never been asked to donate my medical record.
    • I prefer not to answer.
Internet Skills
  1. (1)
    How familiar are you with the following terminology? Please choose a number between 1 and 5 where 1 represents “no understanding” and 5 represents “full understanding” of the term.
    • Advanced search
    • PDF
    • Spyware
    • Wiki
    • Cache
    • Phishing
  2. (2)
    Which of the following best describes your educational background or job field?
    • I have an education in, or work in, the field of computer science, computer engineering or IT
    • I DO NOT have an education in, nor do I work in, the field of computer science, computer engineering or IT.
    • Prefer not to answer.
  3. (3)
    Have you ever heard any of the following techniques? (select all that apply)
    • Differential Privacy
    • End-to-end encryption
    • Public-key cryptography
    • Trusted execution environment (TEE)
    • Pseudo-party computation
    • None of the above
Demographics
  1. (1)
    What year were you born in? (4 digits)
  2. (2)
    To which gender do you most identify? (Select all that may apply)
    • Woman
    • Man
    • Non-binary
    • Prefer not to answer
    • Prefer to self-describe
  3. (3)
    What is the highest level of education you have completed?
    • Less than high school degree (1)
    • High school graduate (high school diploma or equivalent including GED) (2)
    • Some college but no degree (3)
    • Associate’s degree (4)
    • Bachelor’s degree (5)
    • Advanced degree (e.g., Master’s, doctorate) (6)
    • Prefer not to answer (7)
Appendix B Code Book
@@
Data Donation.Data DonationData Donation.
  1. (1)
    Irrelevant response
  2. (2)
    Opt to donate regardless
  3. (3)
    Opt not to donate regardless
  4. (4)
    Require compensation
  5. (5)
    Only donate to a non-profit entity
  6. (6)
    No donation to a for-profit entity
  7. (7)
    Need more information
  8. (8)
    The provided protection is persuasive
  9. (9)
    Generally, the donation scenario looks good
  10. (10)
    Not satisfied with the provided protection
@@
Trustworthy.TrustworthyTrustworthy.
  1. (1)
    Trust without a specific reason
  2. (2)
    Irrelevant response
  3. (3)
    Perfect protection does not exist
  4. (4)
    The recipient entity has no reason not to protect the data
  5. (5)
    The recipient entity has no reason to protect the data
  6. (6)
    A non-profit entity is trustworthy
  7. (7)
    A for-profit entity is not trustworthy
  8. (8)
    Need more information
  9. (9)
    The provided protection is persuasive
  10. (10)
    Generally, the donation scenario looks good
  11. (11)
    Not satisfied with the provided statement

### v3

@@ V Concluding Discussion
Prior work has highlighted the ways in which statements on explaining PETs can fail due to being vague or inaccurate. Our work extends this understanding by showing that even when people clearly understand the protection guarantees, the perceived effectiveness of these guarantees is often limited by non-privacy-related factors. Our participants reported strong pre-existing privacy expectations, even in the absence of explicit privacy statements. We observe that these expectations vary strongly based on the profit model of the entity. For for-profit entities, where baseline privacy expectations are lower, strong privacy statements can effectively elevate these expectations among users. In contrast, the impact on non-profit organizations is minimal because users already hold high privacy expectations for these entities. This disparity in privacy expectations between non-profit and for-profit entities reflects deeper psychological mechanisms that influence how people form judgments about organizational trustworthiness and privacy practices.
@@
The Halo Effect.The Halo EffectThe Halo Effect. Participants’ qualitative responses suggest that their strongly positive privacy expectations for non-profit entities arise from basic moral judgments: non-profits are perceived as more ethical and dedicated to serving the public good because they are not pursuing profit. This findings are consistent with a phenomenon termed “the halo effect”, which describes how an overall positive impression of a particular entity (e.g., the good work that a non-profit does) can lead to ungrounded assumptions about how the organization operates [119, 120]. While prior work in the business literature focuses on how the halo effect influences expectations about business practices such as donation management, our work illustrates that the halo effect also influences people’s expectations about the way non-profits handle data privacy.
However, these positive expectations do not always align with reality. Privacy safeguards are not consistently enforced or widely practiced in non-profit entities; for example, many health organizations rely on broad consent frameworks rather than adopting stricter purpose-restricted consent measures [121]. This gap between expectations and reality can result in privacy expectation violations, a major source of privacy concerns as explained by the theory of contextual integrity [122]. Indeed, a well-known non-profit mental health helpline shared it’s data with a commercial entity, leading to significant public outcry [123]. Future work may seek to explore how to design privacy communications that set appropriate expectations in the presence of such effects, as a prior work has sought to do in other domains (e.g., [124]).
@@
Privacy Washing.Privacy WashingPrivacy Washing. Conversely, our participants had far lower initial expectations of the privacy practices of for-profit entities due to their perceived singular focus on profit (a “horn” effect). As a result, information about data protection guarantees significantly raised their privacy expectations, and in turn, their willingness to donate their data.
This means for-profit entities can gain competitive advantage [125] by transparently implementing and communicating PETs to counter their trust deficit. Yet this dynamic creates vulnerability to “privacy washing” [126, 48], where organizations make vague or exaggerated privacy claims to deceptively raise expectations. A recent incident involving GoodRx exemplifies this problem: despite promising never to share personal health information, it shared sensitive data with third-party advertisers [127]. Such deception ultimately erodes trust and discourages future donations when discovered, deepening the very horn effect that motivated privacy washing.
@@
Opt in Versus Opt out.Opt in Versus Opt outOpt in Versus Opt out. Switching to an “opt-out” model for health data donation might seem appealing to overcome donation hesitancy, particularly for for-profit entities struggling against lower baseline trust. However, such approaches face legal barriers in the U.S. HIPAA [128] sets a high bar for health data sharing, requiring explicit “opt-in” authorization for most research uses.
@@
The Role of Auditing.The Role of AuditingThe Role of Auditing. In theory, auditing could serve as a powerful mechanism to counterbalance both the halo effect and privacy washing by providing objective verification of privacy guarantees, regardless of organizational structure. Auditing should transform abstract guarantees into verifiable technical reality. Yet our findings reveal that auditing has surprisingly limited impact on elevating privacy expectations among participants, with the exception of scenarios involving purpose use restrictions and for-profit entities.
This limited effectiveness arises from participants’ fundamental skepticism about technological infallibility., and shows a disconnect between expert and public perspectives on verification mechanisms, consistent with findings in prior work [32, 67, 34]. While security and privacy experts typically regard auditing as foundational to establishing trust and demonstrating compliance, our participants takes it as merely another layer of protection that could ultimately fail. As one respondent succinctly stated, “They can guarantee privacy all they want; things still get hacked.” This skepticism aligns with emerging research suggesting that many users have developed a resigned attitude toward privacy, doubting that true privacy protection is achievable in practice [73].
@@
Recommendations for Recipient Entities.Recommendations for Recipient EntitiesRecommendations for Recipient Entities. Given participants’ skepticism about privacy protections, recipient entities seeking health data donations should consider several approaches to address these concerns. First, they should recognize that users’ prior experiences with data leakage and exposure to media coverage of breaches create a sense of inevitability about privacy risks. This fatalistic perspective requires more than just technical solutions.
For technical implementations, recipient entities should consider advanced technologies like multi-party computation (MPC) that can minimize breach impacts by distributing sensitive information across multiple entities, ensuring no single point of failure. However, these systems must be engineered carefully, as any failure could damage trust more severely than if simpler methods were compromised.
When communicating privacy guarantees, recipient entities should move beyond explaining basic techniques like E2EE. Our findings show that participants struggle to understand more complex concepts like auditing, despite their effectiveness. Recipient entities should develop communication approaches that better explain these sophisticated cryptographic concepts in accessible ways that demonstrate their benefits.
For-profit entities in particular should recognize their trust disadvantage and focus on transparent implementation and clear communication of robust privacy protections. At the same time, they should avoid “privacy washing” through vague or exaggerated claims, which can erode trust if discovered.
@@
Recommendations for End-Users.Recommendations for End-UsersRecommendations for End-Users. Our research highlights several key points that individuals should consider before donating their health data. First, donors should evaluate recipient entities based on the data protections they explicitly state in policy documents rather than their organizational structure. The halo effect observed in our study demonstrates how implicit assumptions about certain types of entities’ business practices can lead to unfounded assumptions about privacy practices, particularly for non-profit entities.
Second, individuals should avoid assuming privacy protections exist beyond those explicitly stated by the entity requesting data. While entities are legally bound to follow both their explicit promises and applicable privacy laws, the baseline legal protections may be less comprehensive than what users expect for sensitive health data [129].
Finally, individuals should consider engaging privacy advocates before donating sensitive health data. Such consultation can help identify potential privacy risks that might not be immediately obvious and provide a more balanced perspective on the true privacy implications of donation.
@@
Future Work.Future WorkFuture Work. While our study provides valuable insights into privacy expectations and willingness to donate health data, several promising directions remain for future research. First, researchers could expand beyond our four privacy guarantees and two auditing mechanisms to incorporate a more comprehensive range of PETs, with a particular focus on understanding how advanced and complex guarantees can be effectively communicated to non-technical audiences and how they shape privacy expectations and willingness to donate. This could include exploring the interplay between different privacy guarantees and how they are perceived by individuals.
Second, future work should examine these privacy expectations across more diverse demographic groups, including individuals with a strong distrust of data-handling practices. Identifying and understanding these individuals is critical in fostering confidence in PETs and practices.
Third, our study focused on formal medical data donation scenarios, yet health data is increasingly collected through passive means via consumer devices and casual in-app permissions. As prior work emphasizes the importance of modality, timing, and context in privacy information design [125], future research could explore how such factors intersect with biases such as the halo effect to influence privacy expectations, trust, and willingness to share sensitive information.
Finally, the increasingly common partnerships between for-profit and non-profit entities in the medical domain warrant dedicated study. Building on prior work [52] that found deep skepticism toward hybrid entities in COVID-19 contact tracing, future work may seek to further interrogate how such partnerships, which are common in the medical domain, further complicate our understanding of people’s privacy expectations and their sensitivity to statements of data protection.
Acknowledgment
All authors were supported in full or in part by the Max Planck Society during the course of this research. The fourth author additionally acknowledges support from a Google Research Scholar Award “Aligning Technical Data Privacy Protections with User Concerns” and NSF Award # 2429838.